OpenStudy (anonymous):

Find the exact value: tan 47pi/6. PLEASE HELP ME OUT. I HAVE NO IDEA HOW TO FIND THE EXACT VALUE OF THIS.

OpenStudy (goformit100):

@Hero

OpenStudy (noelgreco):

2pi = x pi/6 Can you calculate x ?

OpenStudy (noelgreco):

\[\frac{ x \pi }{ 6 }=2\pi]

OpenStudy (anonymous):

try to use calculator.. :)

OpenStudy (noelgreco):

Yeah. Can you express 2 pi in sixths?

OpenStudy (noelgreco):

The calculator will not give an exact answer.

OpenStudy (noelgreco):

Of course there are tools in technology that provide the answer. The nature of the question suggests that use of the unit circle is required for the solution.

OpenStudy (campbell_st):

this is the same as \[7 \frac{5 \pi}{6}\] which is in the 4nd quadrant and the same as \[2 \pi - \frac{\pi}{6} \] so tan is negative in the 4th quadrant... so just find the exact value of \[\frac{\pi}{6} \] and make it negative.